'Lipin thinks of a number.
She doubles the number and then adds it to the reciprocal of the number.
She gets the answer 3.
Write down and solve an equation to work out the two possible numbers that lipin could be thinking of.
You must show your working.'
any help would be great :)

Answer by sudhanshu_kmr(1152)


let number = x
then 2x + 1/x = 3
=> 2x^2 + 1 = 3x
=> 2x^2 -3x +1 = 0
=>(2x-1)(x-1) =0 , after factorization
so, x = 1/2 , 1